OpenStudy (anonymous):

Find the real solution of x^3+2x^2+5x*10=0 by factoring

OpenStudy (anonymous):

5x*10 or 5x+10?

OpenStudy (anonymous):

\[x ^3 +2x^2 +5x+10\] split it in half and factor so \[x^3 +2x^2 = x^2(x+2) \] \[5x+10= 5(x+2)\] Then it turns into \[(x^2+5)(x+2)\] and you just solve for x